OpenStudy (anonymous):

A motorist travels at a constant speed of 34.0 m/s through a school zone, exceeding the posted speed limit. A policeman waits 7.0 s before giving chase at an acceleration of 3.7 m/s2.

OpenStudy (anonymous):

Find the time required to catch the car from the instant the car passes the policeman.

OpenStudy (anonymous):

Find the distance required for the policeman to overtake the motorist.

OpenStudy (anonymous):

For Car ( insert the values and get an equation) s = ut s = 34 t ------> (1) equation For Policeman ( he starts 7 s later so it is (t-7)) s= ut + 1/2 at^2 ( u = 0) as he strts frm rest \[s = \frac{1}{2} \times 3.7 \times (t-7)^{2}\] \[2s = 3.7(t^{2}-14t+49)\] \[2s = 3.7t^2 -51.8t +181.3\] Now plug in equation {1} in to s so \[2(34t) = 3.7t^{2}-51.8t+181.3\] i think now u can solve and find time.

OpenStudy (anonymous):

did u get it ...

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

ok which one u dont understnd?

OpenStudy (anonymous):

im at 3.7t^2 - 119.8t + 181.3 = 0 now im working on the quadratic equation but I keep getting it wrong please help

OpenStudy (anonymous):

u can either use the calculator or this formula

OpenStudy (anonymous):

i tried the calculator and it keeps saying error

OpenStudy (anonymous):

or insert the values in this http://www.math.com/students/calculators/source/quadratic.htm